OpenStudy (anonymous):

1/2^2/a^2 somehow equals a^2/2^2, I know this, but how? Can someone please explain this?

OpenStudy (anonymous):

Please I'm not looking for answers just help

OpenStudy (perl):

you need parentheses 1 / (2^2 / a^2) = a^2 / 2^2 but 1 / 2^2 / a^2 = 1/(2^2* a^2) by order of operations

OpenStudy (bibby):

division = multiplication by reciprocal. simplify the denominator: \(\huge \frac{1}{\frac{\color{red}{2^2}}{\color{red}{a^2}}}=\frac{1}{}*\frac{a^2}{2^2}\)

OpenStudy (anonymous):

Thanks bibby